\begin{thebibliography}{99}

\bibitem{Belyi}G.~Belyi, Galois extensions of a maximal cyclotomic field. (Russian) Izv. Akad. Nauk SSSR Ser. Mat. 43 (1979), no. 2, 267–276, 479. (English) Math USSR Izv. 14 (1979),. 247-256.

\bibitem{BDH}H.~Braden and L.~Disney-Hogg, Bring's curve: old and new. Eur. J. Math. 10 (2024), no. 1, Paper No. 3, 46 pp.

\bibitem{CV}P.~Clark and J.~Voight, 
Algebraic curves uniformized by congruence subgroups of triangle groups. 
Trans. Amer. Math. Soc. 371 (2019), no. 1, 33–82.

\bibitem{C}M.~Craig, A sextic Diophantine equation.
Austral. Math. Soc. Gaz. 29 (2002), no. 1, 27–29. 

\bibitem{DS}F.~Diamond and J.~Shurman, A first course in modular forms. Graduate Texts in Mathematics, 228. Springer-Verlag, New York, 2005.

\bibitem{Filom}K.~Filom, The Belyi characterization of a class of modular curves. 
J. Th\'eor. Nombres Bordeaux 30 (2018), no. 2, 409–429.

\bibitem{G}E.~Girondo, D.~Torres-Teigell, and J.~Wolfart, 
Fields of definition of uniform dessins on quasiplatonic surfaces. 
Riemann and Klein surfaces, automorphisms, symmetries and moduli spaces, 155–170,
Contemp. Math., 629, Amer. Math. Soc., Providence, RI, 2014.


\bibitem{HM}T.~ Hibino and N.~Murabayashi, Modular equations of hyperelliptic $X_0(N)$ and an application. Acta Arith. 82 (1997), no. 3, 279–291. 


\bibitem{H}K.~ Hulek, Geometry of the Horrocks-Mumford bundle. Algebraic geometry, Bowdoin, 1985 (Brunswick, Maine, 1985), 69–85, Proc. Sympos. Pure Math., 46, Part 2, Amer. Math. Soc., Providence, RI, 1987.

\bibitem{JW}G.~Jones and J.~Wolfart, Dessins d'enfants on Riemann surfaces. Springer Monographs in Mathematics. Springer, Cham, 2016.

\bibitem{KM}N.~Katz and B.~Mazur, Arithmetic moduli of elliptic curves. Annals of Mathematics Studies, 108. Princeton University Press, Princeton, NJ, 1985. 

\bibitem{Kob}N.~ Koblitz, Introduction to elliptic curves and modular forms. Second edition. Graduate Texts in Mathematics, 97. Springer-Verlag, New York, 1993. x+248 pp.

\bibitem{LMFDB}The LMFDB Collaboration, The $L$-functions and modular forms database (https://www.lmfdb.org/).

\bibitem{Mum}D.~Mumford, Tata lectures on theta. I. With the assistance of C. Musili, M. Nori, E. Previato and M. Stillman. Progress in Mathematics, 28. Birkhäuser Boston, Inc., Boston, MA, 1983. xiii+235 pp.

\bibitem{Ribet80}K-A.~Ribet, Twists of modular forms and endomorphisms of abelian varieties.
Math. Ann. 253 (1980), no. 1, 43–62.

\bibitem{Ribet}K.~Ribet, Abelian varieties over $\Q$ and modular forms. Modular curves and abelian varieties, 241–261, Progr. Math., 224, Birkh\"auser, Basel, 2004. 

\bibitem{Se97}J-P.~Serre, Lectures on the Mordell-Weil theorem. Translated from the French and edited by Martin Brown from notes by Michel Waldschmidt. With a foreword by Brown and Serre. Third edition. Aspects of Mathematics. Friedr. Vieweg \& Sohn, Braunschweig, 1997.

\bibitem{Se08}J-P.~ Serre, Topics in Galois theory. Second edition. With notes by Henri Darmon. Research Notes in Mathematics, 1. A K Peters, Ltd., Wellesley, MA, 2008. 

\bibitem{Shimura}G.~ Shimura, On the real points of an arithmetic quotient of a bounded symmetric domain. Math. Ann. 215 (1975), 135–164. 

\bibitem{SM}M.~Shimura, Defining equations of modular curves $X_0(N)$. Tokyo J. Math. 18 (1995), no. 2, 443–456.

\bibitem{SV}J.~Sijsling and J.~Voight, On computing Belyi maps. Num\'ero consacr\'e au trimestre "M\'ethodes arithm\'etiques et applications'', automne 2013, 73–131, Publ. Math. Besancon Algèbre Th\'eorie Nr., 2014/1, Presses Univ. Franche-Comt\'e, Besancon, 2014. 

\bibitem{Sil}J.~ Silverman, The arithmetic of elliptic curves. Second edition. Graduate Texts in Mathematics, 106. Springer, Dordrecht, 2009.

\bibitem{SS}D.~Singerman and R.~Syddall, Belyi uniformization of elliptic curves. Bull. London Math. Soc. 29 (1997), no. 4, 443–451.

\bibitem{Wolfart}J.~Wolfart,
ABC for polynomials, dessins d'enfants and uniformization—a survey. Elementare und analytische Zahlentheorie, 313–345,
Schr. Wiss. Ges. Johann Wolfgang Goethe Univ. Frankfurt am Main, 20, Franz Steiner Verlag Stuttgart, Stuttgart, 2006.

\bibitem{Z}A-K.~Zvonkin, Functional composition is a generalized symmetry, Symmetry: Culture and Science Vol. 2, Nos.1-4, 333-368, 2010. 

\end{thebibliography}

\documentclass[a4paper,11pt]{amsart}
\usepackage{amssymb,mathrsfs}
\usepackage{amsthm}
\usepackage{amsmath}
\usepackage{latexsym}
\usepackage{fancyhdr}
\usepackage{ascmac}
\usepackage{color}
\usepackage[all]{xy}
\usepackage{amscd}

\theoremstyle{plain}
\newtheorem{thm}{Theorem}[section]
\newtheorem{prop}[thm]{Proposition}

\newtheorem{rmk}[thm]{Remark}

\makeatletter

\@addtoreset{equation}{section}
\makeatother

\newcommand{\bQ}{\overline{\mathbb{Q}}}

\newcommand{\C}{\mathbb{C}}
\newcommand{\R}{\mathbb{R}}
\newcommand{\Q}{\mathbb{Q}}

\newcommand{\Z}{\mathbb{Z}}

\newcommand{\F}{\mathbb{F}}

\newcommand{\lra}{\longrightarrow}

\newcommand{\bH}{\mathbb{H}}

\renewcommand{\O}{\mathcal{O}}

\newcommand{\ds}{\displaystyle}

\newcommand{\G}{\Gamma}

\newcommand{\SL}{{\rm SL}}

\newcommand{\e}{\varepsilon}

\newcommand{\bs}{\backslash}

\title[] 
{Computing algebraic Belyi functions on Bring's curve}
\author{Madoka Horie and Takuya Yamauchi}

\keywords{Belyi functions, Bring's curve, Hulek-Craig's curve, modular curves, triangular Shimura curves}
\thanks{}
\subjclass[2010]{}

\address{Madoka Horie \\
Graduate School of Science and Technology, Sophia University\\
 7-1 Kioi-cho, Chiyoda-ku, Tokyo, 102-8554, JAPAN}
\email{horiemaaa@gmail.com}

\address{Takuya Yamauchi \\ 
Mathematical Inst. Tohoku Univ.\\
 6-3,Aoba, Aramaki, Aoba-Ku, Sendai 980-8578, JAPAN}
\email{takuya.yamauchi.c3@tohoku.ac.jp}

\begin{document}

\maketitle

\begin{abstract}
In this paper, we explicitly compute two kinds of algebraic Belyi functions on Bring's curve. 
One is related to a congruence subgroup of $\SL_2(\Z)$ and the other is related to a 
congruence subgroup of the triangle group $\Delta(2,4,5)\subset \SL_2(\R)$. To carry out the computation, we use elliptic cusp forms of weight 2 for the former case and the automorphism group of Bring's curve for the latter case. We also discuss a suitable base 
field (a number field) for describing isomorphisms between Hulek-Craig's curve, Bring's curve, and another algebraic model obtained as a modular curve. 
\end{abstract}

\section{Introduction}
The main result of Belyi \cite{Belyi} shows that a complex projective smooth curve $C$ has a 
model, as an algebraic curve, defined over a number field if and only if there exists a non-constant function $\beta:C\lra \mathbb{P}^1$ 
ramified at most over $\{0,1,\infty\}$. In this paper, we call such a function 
$\beta$ an algebraic Belyi function and more precisely, an algebraic Belyi function 
defined over a number field $K$ if it is defined over $K$. 
Thus, any projective smooth geometrically connected curve over a number field is endowed with 
an algebraic Belyi function. However, finding such a $\beta$ explicitly for a given $C$ is another matter. In fact, 
the explicit computation of algebraic Belyi functions has been studied extensively
(cf. \cite{SS},\cite{SV} among others). 
Belyi's result can also be understood via hyperbolic uniformization and 
there are two types of uniformization:
\begin{enumerate}
\item (Non-compact case) There exists a finite index subgroup $\G$ of $\SL_2(\Z)$ 
and a biholomorphism $X(\G):=\G\bs (\mathbb{H}\cup\mathbb{P}^1(\Q))\simeq C(\C)$ such that 
$\beta$ is interpreted as the natural projection $$X(\G)\lra X(\SL_2(\Z))\stackrel{\sim}{\lra} \mathbb{P}^1(\C)
,\ \G\tau\mapsto \SL_2(\Z)\tau\mapsto \frac{1}{1728}j(\tau)$$  
where $j$ is Klein's $j$-invariant function, with expansion $j(\tau)=q^{-1}+744+196884q+\cdots,\ 
q:=e^{2\pi \sqrt{-1}\tau},\ \tau\in\mathbb{H}:=\{\tau\in\C\ |\ {\rm Im}(\tau)>0\}$. 
\item (Compact case) There exists a finite index subgroup $\Delta$ of 
some triangle group $\Delta(a,b,c)\subset \SL_2(\R)$  
and a biholomorphism $X_\Delta:=\Delta\bs\mathbb{H}\simeq C(\C)$ such that 
$\beta$ is interpreted as the natural projection $$X_\Delta\lra X_{\Delta(a,b,c)}
\stackrel{\sim}{\lra} \mathbb{P}^1(\C)
,\ \Delta\tau\mapsto \Delta(a,b,c)\tau\mapsto j_{\Delta(a,b,c)}(\tau)$$
where $j_{\Delta(a,b,c)}$ is a uniformizer of $X_{\Delta(a,b,c)}$. 
\end{enumerate}
We refer to \cite[p.71, Theorem]{Se97} for the non-compact case and 
\cite[Section 3.3, Theorem 3]{Wolfart} for the compact case.

In this paper, using the above two interpretations, we compute two algebraic Belyi functions on Bring's curve defined by
\begin{equation}\label{Bring}
B:\left\{
\begin{array}{l}
x_0+x_1+x_2+x_3+x_4=0 \\
x^2_0+x^2_1+x^2_2+x^2_3+x^2_4=0 \\
x^3_0+x^3_1+x^3_2+x^3_3+x^3_4=0 
\end{array}\right.
\end{equation}
inside $\mathbb{P}^4$ with homogeneous coordinates $[x_0:x_1:x_2:x_3:x_4]$. 
The symmetric group $S_5$ of degree 5 acts on $B$ by permutation of coordinates and it is known that $S_5$ is 
the full automorphism group of $B$. We refer to \cite[Example 5.5 and Proposition 5.6]{Z} for another aspect of Bring's curve. 
  
A key ingredient in computing the Belyi maps is to relate Bring's curve with a certain modular curve and Hulek-Craig's curve. We briefly introduce 
these curves and the details are given in the corresponding sections.  
Let $X(\G_B)^{{\rm alg}}$ be the algebraic model over $\Q$ of the compactified modular curve $X(\G_B):=\overline{\G_B\bs \mathbb{H}}$ 
which is given by 
\begin{equation}\label{XBalg}
X(\G_B)^{{\rm alg}}:\left\{
\begin{array}{l}
-(x^2+y^2)+2(z^2+w^2)=0 \\
x^3+xy^2+2y^3-4xz^2-8xzw=0
\end{array}\right.
\end{equation}
inside $\mathbb{P}^4$ (see Section \ref{mc}). Here $\G_B:=\Gamma_0(2)\cap \G(5)$ is a congruence subgroup in $\SL_2(\Z)$. 
Let $\beta^{{\rm alg}}_{\G_B}:X(\G_B)^{{\rm alg}}\lra \mathbb{P}^1_\Q$ be an algebraic model over $\Q$ of 
$X(\G_B)\lra X(\SL_2(\Z))\simeq \mathbb{P}^1(\C), \tau\mapsto \frac{1}{1728}j$ 
(see Theorem \ref{belyi1}). 
 
On the other hand, Hulek-Craig's curve $X^{{\rm HC}}$ over $\Q$ is defined by the desingularization of a projective geometrically irreducible singular curve $$F(x_1,y_1,z_1):=x_1(y^5_1+z^5_1)+x^2_1y^2_1z^2_1-x^4_1y_1z_1-2y_1^3z^3_1=0$$
inside $\mathbb{P}^2$ with the coordinates $[x_1:y_1:z_1]$. This curve has studied in \cite{H} and \cite{C}. 
Hulek-Craig's curve $X^{{\rm HC}}$ appears in the context of the Horrocks-Mumford
bundle \cite{H}. Through its symmetries, the latter is also closely related
to the group $\G(5)$.

Then we prove the following result:
\begin{thm}\label{submain}{\rm (}Theorem \ref{isom}{\rm)} Keep the notation as above. Let $\zeta_5$ be a primitive fifth root of unity.  
\begin{enumerate}
\item Bring's curve $B$ is isomorphic over $\Q(\zeta_5)$ to Hulek-Craig's curve $X^{{\rm HC}}$;
\item  Hulek-Craig's curve $X^{{\rm HC}}$ is isomorphic over $\Q$ to $X(\G_B)^{{\rm alg}}$. 
\end{enumerate}
\end{thm}
We note that the first part of Theorem \ref{submain} is due to \cite[Proposition 2.9]{BDH}.
The main results of this paper are Theorem \ref{belyi1} and Theorem \ref{belyi2} and they are summarized as follows: 
\begin{thm}\label{main} Bring curve has two algebraic Belyi functions which are given by the following contexts:  
\begin{enumerate}
\item {\rm (}non-compact type{\rm) }The composition of $B\stackrel{\Q(\zeta_5)}{\simeq}X^{{\rm HC}}
\stackrel{\Q}{\simeq} X(\G_B)^{{\rm alg}}$ and $\beta^{{\rm alg}}_{\G_B}:X(\G_B)^{{\rm alg}}\lra \mathbb{P}^1_\Q$ gives an algebraic Belyi map over $\Q(\zeta_5)$ on $B$. 
\item  {\rm (}compact type{\rm) } The natural quotient $B\lra B/S_5$ gives an algebraic Belyi map over $\Q$ on $B$. 
It is an algebraic model over $\Q$ of the natural quotient map $X_\Delta\lra X_{\Delta(3,4,5)}\simeq \mathbb{P}^1(\C)$ induced by the inclusion of a certain subgroup $\Delta$ of $\Delta(3,4,5)$. 
\end{enumerate}

\end{thm}

Let us explain about some key points for the main results. 
It is known that 
\begin{enumerate}
\item (non-compact case  \cite[p.82-83]{H}) $B(\C)\simeq X(\G_B)$ as a Riemann surface, and  
\item (compact case  \cite[p.73]{CV}) $B(\C)\simeq \Delta\bs \bH$ for some normal congruence subgroup $\Delta$ of $\Delta(3,4,5)$ such that 
$\Delta(3,4,5)/\Delta \simeq  S_5$. Thus,  
${\rm Aut}(B(\C))\simeq {\rm Aut}(\Delta\bs \bH)= \Delta(3,4,5)/\Delta \simeq  S_5$.  
\end{enumerate}

For the non-compact case, we compute an algebraic model of the $j$-function in terms of the modular functions 
on $X(\G_B)$ by regarding $j$ and modular functions with algebraic coordinates. 

For the compact case, we first use the uniformization to conclude that the natural quotient map 
$B(\C)\mapsto B(\C)/{\rm Aut}_{\C}(B(\C))$ is an algebraic Belyi function. 
As a projective smooth complex curve, Bring's curve is the unique curve of genus 4 with $S_5$ as the automorphism group. On the other hands, $S_5$ acts properly on the coordinates of $B$ and 
if we view $B$ as an algebraic curve over $\Q$, we have ${\rm Aut}_\Q(B)\supset S_5$ where 
${\rm Aut}_\Q$ stands for the group of all automorphims defined over $\Q$. However, 
$S_5\subset {\rm Aut}_\Q(B)\subset  {\rm Aut}_\C(B(\C))\simeq S_5$. Thus, ${\rm Aut}_\Q(B)=S_5$. 
Hence, the natural quotient map decent to the quotient map 
$B\lra B/{\rm Aut}_\Q(B)\simeq \mathbb{P}^1_\Q$ which is an algebraic Belyi function defined over $\Q$. 

We will organize this paper as follows. 
In Section \ref{mc}, we study another model of $B$ which is explicitly given as an algebraic model of  the modular curve $X(\G_B)$. 
By using elliptic cusp forms of weight 2 and the theory of canonical curves, we compute 
an explicit defining equation over $\Q$ for $X(\G_B)$. 
In Section \ref{tc}, we will compute the quotient map 
$B\mapsto B/{\rm Aut}_{\Q}(B)$ explicitly. 
Finally, in Section \ref{tab},  we relate our algebraic model of $X(\G_B)$ with Hulek-Craig's curve 
and discuss a number field $K$ such that these curves are 
isomorphic over $K$ to Bring's curve. 
This shows that Bring's curve inherits two interesting algebraic Belyi functions. 

\textbf{Acknowledgments.} We would like to thank Professors K.~Hulek and H.~Braden 
for helpful discussions. We also thank Natalia Amburg for kindly informing us an article \cite{Z}. 
Finally, we would like to give special thanks to the referee, whose suggestions have greatly improved the presentation and readability of this paper.

\section{Bring's curve as a modular curve}\label{mc}
In this section, we compute the Belyi function on Bring's curve $B$ as 
the modular curve $X(\G_B)$ defined below. 
It will be explained in Section \ref{tab} that 
$B$ is isomorphic over $\Q(\zeta_5)$ to an algebraic model over $\Q$ of $X(\G_B)$. Thus, we may work on $X(\G_B)$ to compute the Belyi function. 

We refer to \cite{DS} for the basic facts of elliptic modular forms and modular curves. 
For each positive integer $N$, we denote by $\G(N)\supset \G_0(N)\supset \G_1(N)$ the three kinds of congruence subgroups inside $\SL_2(\Z)$ introduced in \cite[p.13, Definition 1.2.1]{DS}. For such a congruence subgroup $\G$, we define the 
(compact) modular curve $X(\G):=\G\bs (\mathbb{H}\cup\mathbb{P}^1(\Q))$ 
and put 
$X_0(N)=X(\G_0(N))$
 for simplicity.  
We denote by $S_2(\G)$ the space of all cusp forms of weight 2 with respect to $\G$. 
It is known that we have an isomorphism $S_2(\G)\stackrel{\sim}{\lra}H^0(X(\G),\Omega^1_{X(\G)}),\ f(\tau)\mapsto f(\tau)d\tau$ where the right hand side stands for the space of 
all holomorphic 1-forms on $X(\G)$. 
\subsection{Defining equations over $\Q$ of $X(\G_B)$}
Let 
$$\G_B=\Bigg\{
\left(
\begin{array}{cc}
a & b \\
c & d
\end{array}\right)\in \SL_2(\Z)\ \Bigg|\ a\equiv d\equiv 1\ {\rm mod}\ 5,\ c\equiv 0\ {\rm mod}\ 50
\Bigg\}
\subset \G_1(5)\cap \G_0(50).$$
As explained in \cite[Section 2]{BDH}, $B(\C)$ is isomorphic to $X(\G_B)$ as a compact Riemann surface. 
Notice that $\left(
\begin{array}{cc}
5^{-1} & 0 \\
0 & 1
\end{array}\right)\G_B\left(
\begin{array}{cc}
5 & 0 \\
0 & 1
\end{array}\right)=\Gamma_0(2,5):=\G_0(2)\cap \G(5)$ in the notation there 
and also $\G_1(50)\subset \G_B\subset \G_0(50)$.  
Then, the natural inclusion $\G_B\subset \SL_2(\Z)$ yields the quotient map 
\begin{equation}\label{beta1}
\beta_{\G_B}:X(\G_B)\lra X(\SL_2(\Z))\stackrel{\sim}{\lra} \mathbb{P}^1(\C),\ \G_B\tau\mapsto
\SL_2(\Z)\tau\mapsto \frac{1}{1728}j(\tau).
\end{equation}
It is well-known (cf. \cite[Proposition 2.1]{Filom}) that $\beta_{\G_B}$ is a Belyi function ramified only at 
$0,1,\infty$. 
Since each of $X(\G_B)$ and $X(\SL_2(\Z))$ has a structure as a projective smooth algebraic curve defined over $\Q$, the Belyi function $\beta_{\G_B}$ can be seen as an algebraic function 
$\beta^{{\rm alg}}_{\G_B}$ over 
$\Q$. Henceforth, we will explicitly compute $\beta^{{\rm alg}}_{\G_B}$.   

Let us first compute defining equations over $\Q$ of $X(\G_B)$. 
\begin{prop}The modular curve  $X(\G_B)$ has the following  algebraic model over $\Q$ as 
a projective smooth geometrically connected algebraic curve defined over $\Q:$
\begin{equation}\label{X}
X(\G_B)^{{\rm alg}}:\left\{
\begin{array}{l}
-(x^2+y^2)+2(z^2+w^2)=0 \\
x^3+xy^2+2y^3-4xz^2-8xzw=0
\end{array}\right.
\end{equation}
inside $\mathbb{P}^3$ with homogeneous coordinates $[x:y:z:w]$.
\end{prop}
\begin{proof}Under the isomorphism $(\Z/50\Z)^\times\simeq \G_0(50)/\G_1(50)$, 
the subgroup $\G_B/\G_1(50)\subset \G_0(50)/\G_1(50)$ corresponds to 
$\langle \overline{11}:=11\ {\rm mod}\ 50\rangle\simeq \Z/5\Z$. Thus, we have $\G_0(50)/\G_B\simeq 
\Z/4\Z$. It follows from this and \cite{LMFDB} that 
$$S_2(\G_B)=\bigoplus_{\chi\in \widehat{(\Z/50\Z)^\times}\atop \chi^4=1}S_2(\G_0(50),\chi)=
S_2(\G_0(50))\oplus S_2(\G_0(50),\chi_2)$$
where $\chi_2:(\Z/50\Z)^\times=\langle \overline{3}:=3\ {\rm mod}\ 50 \rangle \lra 
\C^\times,\ \overline{3}^i\mapsto (-1)^i$.
We refer to \cite[Section 5.2, p.169]{DS} for the first equality. 
As for generators,  we have $S_2(\G_0(50))=\langle f_1,f_2\rangle$ and $S_2(\G_0(50),\chi_2)=\langle f_3,f_4\rangle$ 
such that their $q$-expansions are given by 
\begin{equation}\label{qexp}
\begin{array}{l}
f_1(\tau):=f_{{\rm 50.2.a.a}}(q)=q - q^2 + q^3 + q^4 - q^6 + 2q^7 - \cdots  \\
f_2(\tau):=f_{{\rm 50.2.a.b}}(q)=q + q^2 - q^3 + q^4 - q^6 - 2q^7 + q^8 -\cdots \\ 
f_3(\tau):=\frac{1}{2}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})+f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))=q - q^4 - q^6 +\cdots \\ 
f_4(\tau):=\frac{1}{2\sqrt{-1}}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})-f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))= q^2 + q^3 - 2 q^7 - q^8 -
\cdots
\end{array}
\end{equation}
where each label $\ast$ of $f_\ast$ corresponds to the corresponding datum on \cite{LMFDB} 
and we can read off the Fourier coefficients up to $q^{99}$ for which it is enough for the computation below.  
Readers may visit the website https://www.lmfdb.org/ModularForm/GL2/Q/holomorphic/ linked in \cite{LMFDB} and 
enter 50 for the level, and 2 for the weight, and either 1 or 2 for the character order. 
Then all labelled forms as above will show up individually.  

Since Bring's curve is non-hyperelliptic and of genus 4, its canonical divisor is very ample.
Hence, by noting $S_2(\G_B)\stackrel{\sim}{\lra}H^0(X(\G_B),\Omega^1_{X(\G_B)}),\ f(\tau)\mapsto f(\tau)d\tau$, we have the canonical embedding:
$$\Phi:X(\G_B)\hookrightarrow \mathbb{P}^3(\C),\ \tau\mapsto 
[x:y:z:w],\ (x,y,z,w):=(f_1(\tau),f_2(\tau),f_3(\tau),f_4(\tau)).$$
By using Petri's theorem as explained in \cite[Section 2.2]{SM} for $X_0(N)$ 
(which also applies to $\G_B$), the image of $\Phi$ is given by the intersection of 
a quadratic equation and a cubic equation in $\mathbb{P}^3$. The claim follows by determining 
the coefficients of those equations by using (\ref{qexp}); see \cite{SM} for computational details.
\end{proof}

\subsection{An algebraic Belyi function on $X(\G_B)^{{\rm alg}}$}
In this subsection, we compute the algebraic function over $\Q$ corresponding to the quotient map $X(\G_B)\lra  
X(\SL_2(\Z))\stackrel{\SL_2(\Z)\tau\mapsto \frac{1}{1728}j(\tau)}{\simeq}  \mathbb{P}^1(\C)$ by using the intermediate modular curve $X_0(50)$ so that the quotient map factors through it:
$$X(\G_B)\stackrel{\pi_1}{\lra}  X_0(50)\stackrel{\pi_2}{\lra} X(\SL_2(\Z))
\stackrel{j\atop \sim}{\lra} \mathbb{P}^1(\C).$$
Here $\pi_1$ and $\pi_2$ are natural projections induced by the inclusion 
$\G_B\subset \G_0(50)\subset \SL_2(\Z)$. 
Let $X_0(50)^{{\rm alg}}$ be the hyperelliptic curve over $\Q$ defined by 
\begin{equation}\label{m50}
t^2=s^6-4s^5-10s^3-4s+1
\end{equation}
such that $X_0(50)^{{\rm alg}}(\C)\simeq X_0(50)$ as a compact Riemann surface (see \cite[p.290]{HM}). 
An algebraic function over $\Q$ for $j\circ \pi_2$ has been computed in \cite[p.290]{HM} 
with (\ref{m50}) is explicitly given by 
$$X_0(50)^{{\rm alg}}_\Q\lra \mathbb{P}^1_\Q,\ (s,t)\mapsto j=h(s,t):=\frac{A(s)+B(s)t}{2s^{25}(s^4-s^3+s^2-s+1)^2}$$
where the polynomials $A(x)$ and $B(x)$ in the one variable $x$ are given in \cite[p.290]{HM}. 
Thus, we have only to compute an algebraic map between $X$ and (\ref{m50}) 
corresponding to $\pi_1$.  
 
Put $T:=y/z=f_1/f_2$ and $S:=z/w=f_3/f_4$ so that $T,S\in \C(X_0(50))$. By using Mathematica version 12.1, we can compute 
a Gr\"oebner basis for the ideal generated by $zT-y,wS-z$ and (\ref{X}) in $\Q[x,y,z,w,T,S]$ such that  
we can find a relation $1 + S^2 + 2 S^3 - 4 T - 4 S^2 T - T^2 - S^2 T^2 + 2 S^3 T^2=0$ in that basis. The curve defined by this equation is isomorphic over $\Q$ to (\ref{m50}) by  
$$(S,T)=\Bigg(\frac{s+1}{s-1},\frac{s^3-s^2+s-1+t}{s(s^2+s+2)}\Bigg),\ 
(s,t)=\Bigg(\frac{S+1}{S-1}, \frac{2 (-2 - 2 S^2 + T + S^2 T + 2 S^3 T)}{(S-1)^3}\Bigg).$$
Thus, we have  proved the following theorem:
\begin{thm}\label{belyi1}
The algebraic function $$\beta^{{\rm alg}}_{\G_B}:X(\G_B)^{{\rm alg}}\lra \mathbb{P}^1_\Q,\ [x:y:z:w]\mapsto 
\frac{1}{1728}h\Bigg(\frac{z+w}{z-w},\frac{2 (w^3 x - 2 w^3 y + w x z^2 - 2 w y z^2 + 2 x z^3)}{y (z - w)^3}  \Bigg)$$ is an algebraic Belyi function defined over $\Q$. It is also an algebraic model over $\Q$ of $X(\G_B)\lra X(\SL_2(\Z))\simeq \mathbb{P}^1(\C),\tau\mapsto \frac{1}{1728}j(\tau)$.
\end{thm}

\begin{rmk} The Belyi function $X(\G_B)\lra X(\SL_2(\Z))\simeq \mathbb{P}^1(\C)$ 
in question is a non-Galois covering of $\mathbb{P}^1(\C)$. Thus, it gives rise to a non-regular dessin on $X(\G_B)$. 
\end{rmk}

\subsection{A decomposition of the Jacobian variety over a number field}\label{decompJ}
In this section, we study a number field $L$ such that the Jacobian variety ${\rm Jac}(X(\G_B)^{{\rm alg}})$ of $X(\G_B)^{{\rm alg}}$ decomposes completely over $L$.
For two abelian varieties $A_1,A_2$ over a number field $K$ and a field extension $M/K$, 
we write $A_1\stackrel{M}{\sim}A_2$ if $A_1$ is isogenous over $M$ to $A_2$. 
For each newform $f$ in $S_2(\G_B)$, we denote by $A_f$ the Shimura's abelian variety over $\Q$ 
of dimension $[\Q_f:\Q]$ where $\Q_f$ is the Hecke field of $f$ (cf. \cite[Section 3]{Ribet}). We note that $A_f$ is $\Q$-simple.  
\begin{prop}\label{decomp} Keep the notation being in Proposition \ref{X}. 
Let $L=\Q(\sqrt{5},\sqrt{5+2\sqrt{5}})$.
Then, it holds that
$${\rm Jac}(X(\G_B)^{{\rm alg}})\stackrel{\Q}{\sim}A_{f_{{\rm 50.2.b.a}}}\times J_0(50)_\Q
\stackrel{\Q}{\sim}A_{f_{{\rm 50.2.b.a}}}\times E_{{\rm 50.a3}}\times E_{{\rm 50.b3}}\stackrel{L}{\sim}
E^4$$
where $E_\ast$ is the elliptic curve over $\Q$ with conductor 50 whose label $\ast$ corresponds to one in \cite{LMFDB} and $E=E_{{\rm 50.a3}}:y^2+xy+y=x^3-x-2$ is an elliptic curve over $\Q$ with $j$-invariant $-\frac{25}{2}$.
\end{prop}
\begin{proof}We refer to 
$$\text{https://www.lmfdb.org/EllipticCurve/2.2.5.1/100.1/a/3}$$
linked from 
 $$\text{https://www.lmfdb.org/ModularForm/GL2/Q/holomorphic/50/2/b/a/}$$
for \cite{LMFDB}. 
The decomposition over $\Q$ follows from \cite[Proposition (2.3)]{Ribet80}. 
It follows from \cite{LMFDB} (see the above websites)  that  
$A_{f_{{\rm 50.2.b.a}}}\stackrel{\Q}{\sim}
{\rm Res}_{\Q(\sqrt{5})/\Q}(E_1)$ where 
$E_1=E_{100.1-{\rm a}3}:y^2+(\e+1)xy+(\e+1)y=x^3+\e x^2,\ \e:=\ds\frac{1+\sqrt{5}}{2}$ with the $j$-invariant $-\frac{25}{2}$. Here ${\rm Res}_{\Q(\sqrt{5})/\Q}$ stands for the Weil-restriction 
from $\Q(\sqrt{5})$ to $\Q$.
Since $E_1$ and $E$ are isomorphic each other over $\bQ$, by \cite[Chapter III, Proposition 3.1]{Sil}, 
there exist $u,r,s,t\in \bQ$ such that $E\stackrel{\sim}{\lra}E_1,\ (x,y)\mapsto 
(u^2x+r,u^3y+u^2 sx+r)$. Solving the system of the algebraic equations, we have
$$s^2+s+\e^2=0,\ r=-\e,\ u=\e^{-2}(1+2s).$$
On the other hand, by  \cite[Chapter III, Proposition 3.1]{Sil} again, $E=E_{{\rm 50.a3}}\stackrel{\Q(\sqrt{5})}{\simeq} E_{{\rm 50.b3}},\ (x,y)\mapsto (u'^2x+r',u'^3y+u'^2 s'x+r')$ with 
$r'=0,\ s' =\ds\frac{1}{10} (-5 + \sqrt{5}),\ t'= \frac{1}{50} (-25 + \sqrt{5}),\ u' = \frac{\sqrt{5}}{5}$. 
Thus, the claim follows. 
\end{proof}

\begin{rmk}\label{decomp-remark} Since $A_{f_{{\rm 50.2.b.a}}}$ is a $\Q$-simple abelian surface, ${\rm Jac}(X)$ 
never be isogenous over $\Q$ to the product of four elliptic curves over $\Q$. 
\end{rmk}

\section{Bring's curve as a triangular (Shimura) curve}\label{tc}
Let $\Delta(2,4,5)\subset \SL_2(\R)$ be the triangle group corresponding to the 
triple $(a,b,c)=(2,4,5)$  (see \cite[Section 2]{CV}) so that 
$F=\Q(\sqrt{2},\sqrt{5})$ and $E=\Q(\sqrt{5})$ in the notation there. 
By \cite[Proposition 5.13]{CV}, there exists an embedding 
$\Delta(2,4,5)\hookrightarrow \SL_2(\O_E)$ where $\O_E$ stands for the ring of inters of $E$. We define $\Delta$ by 
the kernel of $\Delta(2,4,5)\lra \SL_2(\O_E)\lra \SL_2(\O_E/(\sqrt{5}))$ and it yields the 
isomorphism ${\rm Aut}(X_{\Delta})=\Delta(2,4,5)/\Delta\simeq \SL_2(\O_E/(\sqrt{5}))\simeq \SL_2(\F_5)\simeq S_5$. 
To be more precise, $\Delta(2,4,5)/\{\pm 1\}$ can be regarded as the group of the reduced norm 1 
of a maximal order inside the quaternionic algebra $A:=\Big(\frac{-1,\e}{E}\Big)$ over $E$ 
where $\e=\frac{1+\sqrt{5}}{2}$ 
(see \cite[p.53]{CV}). Thus, $X_{\Delta(2,4,5)}=(\Delta(2,4,5)/\{\pm 1\})\bs \bH$ is a Shimura curve 
(of genus zero).  
Therefore, we have a $S_5$-Galois covering $X_\Delta\lra X_{\Delta(2,4,5)}$ between 
triangular Shimura curves which is a Belyi function by construction.  
The quotient map $B\lra B/{\rm Aut}_\Q(B)=B/S_5\simeq \mathbb{P}^1_\Q$ is an algebraic 
Belyi function over $\Q$ since its base change to $\C$ is corresponding to the quotient map $X_{\Delta(2,4,5)}\lra X_\Delta$. 
Thus, we have only to find out a generator of the function field of $B/S_5$. 
Let $\rho:B\lra B$ and $\kappa:B\lra B$ be the automorphism corresponding to 
$(12345),(12)\in S_5$ respectively. Notice that $\rho,\kappa$ generate $S_5$.  
Put $T=(x_1x_2x_3x_4)/x^4_5$ and 
\begin{equation}\label{t}
t:=\sum_{i=1}^5\rho^i (T)+\sum_{i=1}^5 \kappa \rho^i (T).
\end{equation}
By direct computation, we can check $t$ is invariant under the action of $S_5$. 
\begin{thm}\label{belyi2}Keep the notation being as above. 
The rational function $t$ is a generator of $\Q(B/S_5)$. It gives rise to an algebraic Belyi function 
over $\Q$:
$$B\lra \mathbb{P}^1,\ [x_1:x_2,x_3:x_4:x_5]\mapsto t.$$ 
\end{thm}
\begin{proof}
Put $v:=x_1/x_0$ for simplicity. 
By using Grobner basis, we have $t=\ds\frac{-2g(v)}{v^4 (1 + v)^4 (1 + v^2)^4}$ where 
$$g(v)=1 + 5 v + 15 v^2 + 35 v^3 + 65 v^4 + 101 v^5 + 135 v^6 + 
   155 v^7 + 165 v^8 + 165 v^9 + 161 v^{10} $$
   $$+ 165 v^{11} +165 v^{12} + 
   155 v^{13} + 135 v^{14} + 101 v^{15} + 65 v^{16} + 35 v^{17} + 15 v^{18} + 
   5 v^{19} + v^{20}.$$
   For a generic $t\in \Q$, we have the number of $v$ satisfying the above equation is 20. 
Further, for each $v$, we have $6$ pairs of $(x_2/x_5,x_3/x_5)$ modulo the equality 
$\ds\sum_{i=1}^5 x_i=0$. 
Thus, $[\Q(B):\Q(t)]=20\times 6=120$. 
On the other hand, we observe that $[\Q(B):\Q(B/S_5)]=[\C(B):\C(B/S_5)]=|\Delta(2,3,5)/\Delta|=|S_5|=120$ and 
$\Q(t)\subset \Q(B/S_5)\subset  \Q(B)$. Thus, the claim follows. 
\end{proof}
\begin{rmk}Since any {\rm(}compact{\rm)} Shimura curve does not posses any real points {\rm(}cf. \cite{Shimura}{\rm)},  
$t$ can not generate the function field of any rational model over $E=\Q(\sqrt{5})$ of $X_{\Delta(2,4,5)}$. 
\end{rmk}

\section{A relation between two algebraic models}\label{tab}
In this section, we study the field of definition of isomorphisms between $B$, $X(\G_B)^{{\rm alg}}$ and the 
Hulek-Craig curve. 
It is known that Bring's curve $B$ is isomorphic to the algebraic model $X(\G_B)^{{\rm alg}}$ of $X(\G_B)$ 
over $\bQ$. We would like to discuss a number field $K$ such that $B$ is isomorphic over $K$ to $X(\G_B)^{{\rm alg}}$.  Obviously, $K$ cannot be $\Q$ since $B(\Q)$ is empty because of the second quadratic equation in (\ref{Bring}) 
while obviously $[1:1:1:0]$ is a $\Q$-rational point on $X(\G_B)^{{\rm alg}}$.  
For another reason, it is known that ${\rm Jac}(B)\stackrel{\Q}{\sim}
E^4$ (cf. \cite[p.87, Exercise-b)]{Se08}) where $E$ is given in Proposition \ref{decomp}.  
Thus, the field $K$ cannot be $\Q$ since ${\rm Jac}(B)$ is not isogenous over $\Q$ to 
${\rm Jac}(X)$ (see also Remark \ref{decomp-remark}). We should note that ${\rm Jac}(X)$ is isogenous over 
$L$ to ${\rm Jac}(B)$, but it is not necessarily true that $X$ is isomorphic over $L$ to $B$. 
Thus, we need more careful study for our purpose.

However, fortunately, it is known from \cite[Proposition 2.9 and its proof]{BDH} that $B$ is isomorphic over $\Q(\zeta_5)$ to Hulek-Craig's curve $X^{{\rm HC}}$ which is 
given by the desingularization of a projective geometrically irreducible singular curve
$$F(x_1,y_1,z_1):=x_1(y^5_1+z^5_1)+x^2_1y^2_1z^2_1-x^4_1y_1z_1-2y_1^3z^3_1=0$$
inside $\mathbb{P}^2$ with the coordinates $[x_1:y_1:z_1]$. This curve has studied in \cite{H} and \cite{C}.  
As explained in \cite[p.27]{C} that Hulek-Craig's curve has 
the parametrization in terms of modular forms:
$$x_1=\sum_{n\in \Z}(-q)^{(5n)^2},\ y_1=\sum_{n\in \Z}(-q)^{(5n+1)^2},\ 
z_1=\sum_{n\in \Z}(-q)^{(5n+2)^2},\ q=e^{2\pi \sqrt{-1}\tau}$$
where we slightly modified $q$ with $-q$ from Craig's notation. 
In terms of the Jacobi's theta function 
$\theta_{a,b}(z,\tau)$, $(z,\tau)\in \C\times \mathbb{H}$ with the rational character $(a,b)\in \Q^2$ (see \cite[p.10]{Mum}), we see that 
$$x_1=\theta_{0,0}(0,25(2\tau+1)),\ y_1=\theta_{\frac{1}{5},0}(0,25(2\tau+1)),\ 
z_1=\theta_{\frac{2}{5},0}(0,25(2\tau+1))$$
and thus, $x_1,y_1,z_1$ are modular forms of weight $\frac{1}{2}$ with respect to 
the double covering of $\G(5)\cap \G_1(4)$ (see \cite[Chapter IV]{Kob}).
It is known (cf. \cite[Lemma 2.6]{BDH}) that 
$$v_1=\frac{(Y^3-X)dX}{F_Y(X,Y,1)},\ 
v_2=\frac{(Y^2 X-1)dX}{F_Y(X,Y,1)},\ 
v_3=\frac{(Y-X^2)dX}{F_Y(X,Y,1)},\ 
v_4=\frac{Y(X^2-Y)dX}{F_Y(X,Y,1)}$$
where $X=x_1/z_1,\ Y=y_1/z_1$, and $F_Y(X,Y,1):=\frac{\partial F(X,Y,1)}{\partial Y}$. 
Then, in terms of $q$-expansion, we have 
$$(v_1,v_2,v_3,v_4)=
(-q^3 + q^8 + 4 q^{13}\cdots ,q^2 - 2 q^7 - q^{12}\cdots, -q^4 + 2 q^9 + 2 q^{14}\cdots, -q + q^6 + 3 q^{11}\cdots  )dq$$
$$=(f_1,f_2,f_3,f_4)A(dq),\ 
A=\left(\begin{array}{cccc}
-\frac{1}{4} & -\frac{1}{4} & -\frac{1}{4} & -\frac{1}{4}  \\   
\frac{1}{4}  & \frac{1}{4} & -\frac{1}{4} & -\frac{1}{4}  \\   
              0& 0 & \frac{1}{2} & -\frac{1}{2}  \\   
-\frac{1}{2} & \frac{1}{2} & 0   & 0
\end{array}
\right)
$$
Put $w_i:=F_Y(X,Y,1)\ds\frac{v_i}{dX}$ for $i=1,2,3,4$ and let us regard $[f_1 dq:f_2 dq:f_3 dq:f_4 dq]$ with 
the coordinates $[x:y:z:w]$ of $X$ in view of the theory of canonical curves. 
Then, $[x:y:z:w]=[w_1:w_2:w_3:w_4]A^{-1}$. Using this, we have a birational map $\psi$ from 
the singular model of $X^{{\rm HC}}$ defined by $F(x,y,z)$ to $X(\G_B)^{{\rm alg}}$ given by 
$\psi([x_1:y_1:z_1])=$
\begin{equation}\label{HC}
\psi([x_1:y_1:z_1])=[x:y:z:w]
\end{equation}
where 
$$\left\{\begin{array}{l}
x=(z_1-y_1) (x^2_1+x_1 y_1+y^2_1+x_1 z_1+z^2_1),\\
y=(y_1-z_1) (-x^2_1+x_1 y_1+y^2_1+x_1 z_1+2 y_1 z_1+z^2_1),\\
z=(y_1+z_1) (-x^2_1+y_1 z_1),\\
w=x_1 y^2_1-y^3_1+x_1 z^2_1-z^3_1.
\end{array}\right.
$$
Then, it yields an isomorphism over $\Q$ from $X^{{\rm HC}}$ to $X(\G_B)^{{\rm alg}}$. 
This result is not obvious from the construction explained in \cite[p.82-83]{H}.  
Thus, we have checked the following result:
\begin{thm}\label{isom}Recall that the projective smooth model $X(\G_B)^{{\rm alg}}$ over $\Q$ of the modular curve $X(\G_B)$ from (\ref{X}).
It holds that 
\begin{enumerate}
\item Bring's curve $B$ is isomorphic over $\Q(\zeta_5)$ to $X^{{\rm HC}}$ whose isomorphism is given by \cite[Proposition 2.9]{BDH}, and  
\item Hulek-Craig's curve $X^{{\rm HC}}$ is isomorphic over $\Q$ to $X(\G_B)^{{\rm alg}}$ whose isomorphism is given by {\rm (\ref{HC})}. In particular, Hulek-Craig's curve is isomorphic over $\Q$ to 
a canonical model of $X(\G_B)$ {\rm(}see \cite{KM} for canonical models of modular curves{\rm)}.
\end{enumerate} 
\end{thm}
Thus, $B$ inherits two algebraic Belyi functions 
such that one is defined  over $\Q(\zeta_5)$ and non-regular (not a Galois covering) while 
the other is defined over $\Q$ and regular (a Galois covering) (see \cite{G} for the 
definition of regular Belyi functions).

\end{document}